### v3

# Wasserstein Barycenter Soft Actor-Critic

Abstract
Deep off-policy actor-critic algorithms have emerged as the leading framework for reinforcement learning in continuous control domains. However, most of these algorithms suffer from poor sample efficiency, especially in environments with sparse rewards. In this paper, we take a step towards addressing this issue by providing a principled directed exploration strategy. We propose Wasserstein Barycenter Soft Actor-Critic (WBSAC) algorithm, which benefits from a pessimistic actor for temporal difference learning and an optimistic actor to promote exploration. This is achieved by using the Wasserstein barycenter of the pessimistic and optimistic policies as the exploration policy and adjusting the degree of exploration throughout the learning process. We compare WBSAC with state-of-the-art off-policy actor-critic algorithms and show that WBSAC is more sample-efficient on MuJoCo continuous control tasks.
1 Introduction
Deep reinforcement learning (DRL) has found applications across diverse fields, including autonomous driving, robotics, healthcare, finance Deng et al. (2016), and gaming Kiran et al. (2021); Hambly et al. (2023). Among the various DRL methodologies, actor-critic frameworks Silver et al. (2014); Lillicrap et al. (2016); Fujimoto et al. (2018) have demonstrated significant success in continuous control tasks and have gained widespread adoption within the control systems and robotics fields. Despite their successes, these methods still face the challenge of high sample complexity, which results from maximizing the lower bound of the Q-function to prevent overestimation and reliance on basic exploration techniques. In practice, the state-of-the-art actor-critic algorithms, such as deep deterministic policy gradient (DDPG) Lillicrap et al. (2016) and its successor, twin delayed DDPG (TD3) Fujimoto et al. (2018), inject symmetric noise directly into the action space to encourage exploration. In contrast, soft actor-critic (SAC) encourages stochasticity through entropy regularization. However, these heuristic exploration approaches are not sample efficient and can limit performance, particularly in sparse-reward or high-dimensional environments Hao et al. (2023).
Moving beyond simple noise injection or basic entropy regularization towards more directed or uncertainty-aware exploration strategies, several approaches have been proposed for actor-critic methods to enhance the performance without causing overestimation and instability Zheng12 et al. (2018); Lee et al. (2021); Januszewski et al. (2021); Likmeta et al. (2023); Ciosek et al. (2019). Most of these efforts employ the principle of optimistic in the face of uncertainty (OFU) Ziebart (2010) and quantify epistemic uncertainty as a signal for exploring promising areas. These works rely on a conservative actor that outputs a pessimistic policy for temporal difference learning and consider an optimistic policy to facilitate exploration. To prevent instability and avoid sub-optimal behavior, the optimistic policy is typically regularized by a Kullback–Leibler (KL) divergence constraint that keeps it close to the pessimistic policy. The other existing works use multiple actors Zhang and Yao (2019); Lyu et al. (2022); Li et al. (2023a); Kumar et al. (2018); Ren et al. (2021); Nauman and Cygan (2025); Yang et al. (2024); Whitney et al. (2021); Li et al. (2023b) to more explicitly decouple the task of temporal difference learning from exploration. While these works have shown successful performance in improving the sample efficiency, establishing a balance between exploration and exploitation is still needed to avoid divergence and suboptimal performance.
Figure 1: WBSAC uses Wasserstein barycenter of optimistic and pessimistic policies as the exploration policy.
To mitigate the aforementioned limitations, we present Wasserstein barycenter soft actor-critic (WBSAC), which uses two distinct actors with different objectives. The optimistic actor is trained to maximize the Q-function upper bound, while the pessimistic actor maximizes a Q-function lower bound. Unlike many approaches that rely on KL-divergence constraints to keep exploration close to a target policy, WBSAC forms the exploration policy as the Wasserstein barycenter of distinct pessimistic and optimistic policies, as shown in Figure 1. This allows for a controlled and geometrically meaningful balance, where the influence of the optimistic policy gradually increases, rather than being strictly constrained. This adaptive balance improves exploration capability and performance without introducing overestimation bias. Using Wasserstein barycenter for blending optimistic and pessimistic strategies and controlling the degree of exploration is the novelty of our work. The main contributions of this paper are:
  • We present WBSAC, an off-policy actor–critic framework with dual actors to enhance exploration capability and performance. WBSAC makes the fusion of pessimistic and optimistic strategies possible in a unique way and allows for a controlled, progressive shift in exploration strategy, starting from conservative behavior towards an optimistic strategy.
  • We theoretically show that WBSAC enjoys high differential entropy, which leads to more effective learning and improved state-action coverage.
  • We empirically demonstrate that WBSAC outperforms state-of-the-art baselines in both sample efficiency and overall performance on several Mujoco tasks and sparse reward case studies.
2 Related work
Significant research has been devoted to optimistic exploration strategies to improve sample efficiency by using Q-value upper bounds. These approaches include analytically derived transformations of pessimistic policies or employing separate exploration-focused actor networks. Directed exploration approaches that use an optimistic policy for exploration and a conservative policy for exploitation, provide different strategies for blending pessimism with optimism Zheng12 et al. (2018); Lee et al. (2021); Januszewski et al. (2021); Likmeta et al. (2023); Ciosek et al. (2019); Moskovitz et al. (2021). For instance, optimistic actor-critic (OAC) Ciosek et al. (2019) models the optimistic policy by combining conservative policy with the linear approximation of the Q-function upper bound, and tactical optimism and pessimism (TOP) Moskovitz et al. (2021) determines the appropriate level of optimism/pessimism during training by framing this selection process as a multi-armed bandit problem. In these works, both optimistic and pessimistic policies are extracted from a conservative actor, and by the KL divergence constraint, the exploration policy stays close to the pessimistic policy. WBSAC is different from these works in both modeling the optimistic policy with a separate actor and the exploration policy. Furthermore, unlike WBSAC, both TOP and Wasserstein actor-critic (WAC) Likmeta et al. (2023) use a distributional critic.
Approaches with multiple actors are similar to WBSAC Nauman and Cygan (2025); Lyu et al. (2022); Li et al. (2023a); Pan et al. (2020); Lyu et al. (2023); Chen et al. (2024); Nauman et al. (2024). WBSAC distinguishes itself from decoupled actor-critic (DAC) Nauman and Cygan (2025), which optimizes the optimistic actor towards a KL-regularized Q-value upper-bound by introducing the Wasserstein barycenter of the policies as the exploration policy. Unlike WBSAC, which concentrates on stochastic policies, DARC Lyu et al. (2022) implements the dual actor setup for deterministic policies. It promotes exploration by selecting the action associated with the higher Q-value from two separate actors and reduces estimation bias through critic regularization. However, basing action selection on the maximal Q-value can amplify errors arising from inaccurate value estimates. Bigger, regularized, optimistic (BRO) Nauman et al. (2024) uses dual policy optimistic exploration and non-pessimistic quantile Q-value approximation to make a balance between exploration and exploitation. Multi-actor mechanism (MAM) Li et al. (2023a) proposes a multi-actor framework with multiple action choices within a single state to optimize policy selection.
Optimal transport has been integrated to RL in several studies Baheri (2023a); Queeney et al. (2023); Metelli et al. (2019); Shahrooei and Baheri (2024); Baheri et al. (2025, 2024); Baheri (2023b). For instance, Metelli et al. (2019) proposes a novel approach called Wasserstein Q-learning (WQL), which considers Bayesian posterior Q and target distributions and applies Wasserstein barycenters to model and propagate uncertainty. Their method demonstrated improved exploration and faster learning in tabular domains compared to classic RL algorithms. Likmeta et al. (2023) extends this idea to actor-critic algorithms and suggests a distributional critic, which minimizes the Wasserstein distance between Q and target distributions. Baheri et al. (2025) introduces an actor-critic algorithm called Wasserstein adaptive value estimation (WAVE), which enhances stability through an adaptive critic regularization term based on the Wasserstein distance between consecutive Q-distributions.
3 Preliminaries
Markov Decision Processes and Q-functions. A Markov Decision Process (MDP) provides a formal framework for sequential decision-making problems in RL. An MDP is defined by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R,\gamma)$, where $\mathcal{S}$ is the set of states, $\mathcal{A}$ is the set of actions, $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ is the state transition probability function (with $\Delta(\mathcal{S})$ being the set of probability distributions over $\mathcal{S}$), $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the reward function, and $\gamma\in[0,1)$ is the discount factor. The behavior of an agent is described by a policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$, which maps states to a probability distribution over actions. The performance of a policy is often evaluated using value functions. The action-value function (or Q-function) for a policy $\pi$, denoted $Q^{\pi}(s,a)$, is the expected discounted return starting from state $s$, taking action $a$, and subsequently following policy $\pi$: $Q^{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right]$. The primary goal in many RL problems is to find an optimal policy $\pi^{*}$ that maximizes the expected return. This optimal policy satisfies $\pi^{*}(s)=\arg\max_{a\in\mathcal{A}}Q^{*}(s,a)$ for all $s\in\mathcal{S}$, where $Q^{*}(s,a)$ is the optimal action-value function.
Soft Actor-Critic. Soft actor-critic (SAC) Haarnoja et al. (2018) is an off-policy actor-critic algorithm based on the maximum entropy RL framework Ziebart et al. (2008). This framework encourages exploration by augmenting the expected return objective with an entropy term for the policy. SAC employs a stochastic policy $\pi_{\phi}$, typically a neural network parameterized by $\phi$. To mitigate overestimation bias, it utilizes two separate Q-function networks, $Q_{\theta_{1}}$ and $Q_{\theta_{2}}$ (parameterized by $\theta_{1}$ and $\theta_{2}$), and uses the minimum of their predictions for policy improvement and target calculation. Correspondingly, two target Q-networks, $\hat{Q}_{\hat{\theta}_{1}}$ and $\hat{Q}_{\hat{\theta}_{2}}$, are maintained and updated via polyak averaging of the Q-network parameters $\hat{\theta}_{i}\leftarrow\tau\theta_{i}+(1-\tau)\hat{\theta}_{i}$, where $\tau\ll 1$. The Q-functions $Q_{\theta_{i}}$ are trained to minimize the soft Bellman residual. For each critic $Q_{\theta_{i}}$, the loss function is:
$\mathcal{L}_{Q}(\theta_{i})=\mathbb{E}_{(s_{t},a_{t},r_{t},s_{t+1})\sim\mathcal{D}}\left[\left(Q_{\theta_{i}}(s_{t},a_{t})-y_{t}\right)^{2}\right],\quad i\in\{1,2\}$ (1)
where the target value $y_{t}=r_{t}+\gamma\left(\min_{j=1,2}\hat{Q}_{\hat{\theta}_{j}}(s_{t+1},a^{\prime}_{t+1})-\alpha\log\pi_{\phi}(a^{\prime}_{t+1}|s_{t+1})\right),\quad a^{\prime}_{t+1}\sim\pi_{\phi}(\cdot|s_{t+1})$ where $\alpha$ is the entropy temperature coefficient, and $\mathcal{D}$ is the replay buffer. The policy $\pi_{\phi}$ is optimized to maximize the expected future return, including the entropy term. This is achieved by minimizing the following loss:
$\mathcal{L}_{\pi}(\phi)=\mathbb{E}_{s_{t}\sim\mathcal{D},\,a_{t}\sim\pi_{\phi}}\Bigl{[}\alpha\,\log\pi_{\phi}(a_{t}\mid s_{t})\;-\;\min_{j=1,2}Q_{\theta_{j}}(s_{t},a_{t})\Bigr{]}$ (2)
The entropy temperature $\alpha$ can be automatically tuned to balance the reward and entropy terms. The loss function for $\alpha$ is:
$\mathcal{L}_{\alpha}(\alpha)=\mathbb{E}_{s_{t}\sim\mathcal{D},a_{t}\sim\pi_{\phi}(\cdot|s_{t})}\left[-\alpha(\log\pi_{\phi}(a_{t}|s_{t})+\mathcal{H}_{0})\right]$ (3)
where $\mathcal{H}_{0}$ is the target entropy.
Wasserstein Barycenter. The Wasserstein distance is a key metric from optimal transport theory to quantify the difference between probability distributions. For a metric space $(\mathcal{X},d)$ equipped with the Borel $\sigma$-algebra $\mathcal{B}(\mathcal{X})$, let $\mu$ and $\nu$ be probability measures defined on this space. Given a proper cost function $d(x,y)$, the $p$-Wasserstein distance Villani (2009) between $\mu$ and $\nu$ is defined as:
$W_{p}(\mu,\nu)=\left(\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathcal{X}\times\mathcal{X}}d(x,y)^{p}\,d\pi(x,y)\right)^{\frac{1}{p}},\quad p\geq 1$ (4)
where $\Pi(\mu,\nu)$ is the set of all couplings (joint probability measures) on the product space $\mathcal{X}\times\mathcal{X}$ with marginals $\mu$ and $\nu$. In this work, we take the cost function to be the Euclidean norm, $d(x,y)=\|x-y\|_{2}$, and set $p=2$. Building upon this notion of distance, Wasserstein barycenters are widely used for capturing the average of probability distributions in a geometrically meaningful way. Given a set of probability measures $\{\mu_{i}\}_{i=1}^{n}$ defined on a metric space $(\mathcal{X},d)$, the Wasserstein barycenter Agueh and Carlier (2011) represents a central probability distribution that minimizes the weighted sum of Wasserstein distances to each distribution in the set. Precisely, given weights $\{\xi_{i}\}_{i=1}^{n}$ satisfying $\sum_{i=1}^{n}\xi_{i}=1$ and $\xi_{i}\geq 0$, the Wasserstein barycenter $\bar{\mu}$ is defined as:
$\bar{\mu}=\arg\min_{\mu}\sum_{i=1}^{n}\xi_{i}W_{2}(\mu,\mu_{i})^{2}$ (5)
4 Wasserstein barycenter soft actor-critic
Our WBSAC algorithm maintains the core components of SAC, including dual critic networks and their targets for Q-value estimation, but distinguishes itself by employing two specialized actor networks and dynamically blending their outputs via Wasserstein barycenter to form the exploration policy.
Algorithm 1 Wasserstein Barycenter Soft Actor-Critic (WBSAC)
Input: exploration schedule parameter $\lambda$, target update rate $\tau$, learning rates $\alpha_{\theta}$, $\alpha_{\pi_{o}}$, $\alpha_{\pi_{p}}$, $\alpha_{\alpha}$, uncertainty bonus parameter $\beta_{o}$
Initialize critic networks $Q_{\theta_{1}}$, $Q_{\theta_{2}}$ with random parameters $\theta_{1}$, $\theta_{2}$
Initialize target networks $\theta_{1}^{{}^{\prime}}\leftarrow\theta_{1}$, $\theta_{2}^{{}^{\prime}}\leftarrow\theta_{2}$
Initialize actor networks with random parameters $\phi$ and $\varphi$
Initialize replay buffer $\mathcal{D}$
1:for each iteration do
2:  for each environment step do
3:   Obtain $\pi_{p}(\cdot|s_{t})$ from pessimistic actor network
4:   Obtain $\pi_{o}(\cdot|s_{t})$ from optimistic actor network
5:   Compute exploration policy $\pi_{e}(\cdot|s_{t})$ according to Eq. 8 with weights $\xi_{p},\xi_{o}$
6:   $a_{t}\sim\pi_{e}(a_{t}|s_{t})$ $\triangleright$ Sample action from the exploration policy
7:   $s_{t+1}\sim p(s_{t+1}|s_{t},a_{t})$ $\triangleright$ Sample transition from the environment
8:   $\mathcal{D}\leftarrow\mathcal{D}\cup\{(s_{t},a_{t},r(s_{t},a_{t}),s_{t+1})\}$ $\triangleright$ Store the transition in the replay buffer
9:   Update Wasserstein barycenter weights $\xi_{o}$ and $\xi_{p}$
10:  for each gradient step do
11:   Sample mini-batch of NN transitions $(s_{t},a_{t},r(s_{t},a_{t}),s_{t+1})$ from $\mathcal{D}$
12:   $\theta_{i}\leftarrow\theta_{i}-\alpha_{\theta}\nabla_{\theta_{i}}\mathcal{L}_{Q}(\theta_{i})\quad\text{for}\quad i\in\{1,2\}$ $\triangleright$ Update critic parameters
13:   $\varphi\leftarrow\varphi-\alpha_{\pi_{o}}\nabla_{\varphi}\mathcal{L}_{\pi_{o}}(\varphi)$ $\triangleright$ Update optimistic policy parameters
14:   $\phi\leftarrow\phi-\alpha_{\pi_{p}}\nabla_{\phi}\mathcal{L}_{\pi_{p}}(\phi)$ $\triangleright$ Update pessimistic policy parameters
15:   $\alpha\leftarrow\alpha-\alpha_{\alpha}\nabla_{\alpha}\mathcal{L}_{\alpha}(\alpha)$ $\triangleright$ Update entropy temperature
16:   $\theta^{\prime}_{i}\leftarrow\tau\theta_{i}+(1-\tau)\theta^{\prime}_{i},\quad\text{for }i=1,2$ $\triangleright$ Update target critic parameters
Output: Optimized parameters $\theta_{1}$, $\theta_{2}$, $\phi$, $\varphi$, $\alpha$
Pessimistic and Optimistic Policies. We first define a pessimistic actor, parameterized by $\phi$, which relies on the lower bound of the Q-function estimates to ensure robust performance. Its objective is formulated as:
$\mathcal{L}_{\pi_{p}}(\phi)=\mathbb{E}_{s\sim\mathcal{D},\,a_{t}\sim\pi_{p}}\left[\alpha\log\pi_{p}(a_{t}|s_{t})-\min_{i=1,2}Q_{\theta_{i}}(s_{t},a_{t})\right]$ (6)
This actor outputs a pessimistic policy denoted by $\pi_{p}\bigl{(}\cdot\mid s_{t}\bigr{)}$. Additionally, WBSAC incorporates an optimistic actor, parameterized by $\varphi$, which explores actions with potentially high long-term rewards using discrepancies between the Q-value predictions of two critic networks. High divergence in critic estimates often indicates regions of the action space with greater uncertainty, where exploration may uncover high-reward actions. This optimistic actor targets these regions by maximizing an objective that balances the average Q-value with a standard deviation term:
$\mathcal{L}_{\pi_{o}}(\varphi)=\mathbb{E}_{s\sim\mathcal{D},\,a_{t}\sim\pi_{o}(\cdot|s_{t};\varphi)}\left[-\left(\mu_{Q}+\beta_{o}\cdot\sigma_{Q}\right)\right]$ (7)
where $\mu_{Q}=\frac{Q_{\theta_{1}}(s_{t},a_{t})+Q_{\theta_{2}}(s_{t},a_{t})}{2}\quad$ and $\sigma_{Q}=\sqrt{\frac{1}{2}\sum_{i=1}^{2}(Q_{i}-\mu_{Q})^{2}}$ and $\beta_{o}$ is a positive hyperparameter, which controls the uncertainty bonus. This actor outputs an optimistic policy specified by $\pi_{o}\bigl{(}\cdot\mid s_{t}\bigr{)}$.
Wasserstein Barycentric Exploration Policy. Since SAC employs stochastic policies by outputting probability distributions over the action space, we propose integrating the pessimistic and optimistic policies through the Wasserstein barycenter. Specifically, we define the exploration policy $\pi_{e}$ as the distribution that minimizes the weighted sum of Wasserstein distances to $\pi_{p}$ and $\pi_{o}$:
$\pi_{e}=\arg\min_{\pi}\left(\xi_{p}\,W_{2}(\pi,\pi_{p})^{2}+\xi_{o}\,W_{2}(\pi,\pi_{o})^{2}\right)$ (8)
where $\xi_{p}$ and $\xi_{o}$ are non-negative respective weights of pessimistic and optimistic policies, which satisfy $\xi_{p}+\xi_{o}=1$. By tuning $\xi_{p}$ and $\xi_{o}$, we ensure a controlled balance between exploitation and exploration.
Exploration Degree Adjustment. To ensure stable learning, WBSAC initially biases $\pi_{e}$ towards the pessimistic policy $\pi_{p}$. This is achieved by scheduling the weight $\xi_{o}$ to start at zero and grow linearly as training progresses. The weight scheduling strategy is considered $\xi_{o}=\min\left(1,\,\lambda\cdot\frac{t}{T}\right)$, where $\lambda$ is the exploration schedule parameter, $t$ denotes the current environment timestep, and $T$ represents the total environment time steps.
Closed-Form Barycenter for Factorized Gaussian Policies. For Gaussian policies, the barycenter has a closed-form solution, which decreases computation cost while preserving theoretical validity. Assuming both the pessimistic policy $\pi_{p}(\cdot|s)=\mathcal{N}(\mu_{p}(s),\Sigma_{p}(s))$ and the optimistic policy $\pi_{o}(\cdot|s)=\mathcal{N}(\mu_{o}(s),\Sigma_{o}(s))$) are Gaussian, their $W_{2}$-barycenter $\pi_{e}(\cdot|s)=\mathcal{N}(\mu_{e}(s),\Sigma_{e}(s))$ is also Gaussian. Its mean $\mu_{e}(s)$ and covariance $\Sigma_{e}(s)$ are efficiently computed in closed form:
$\mu_{e}(s)$ $=\xi_{p}\mu_{p}(s)+\xi_{o}\mu_{o}(s)$ (9)
$\Sigma_{e}(s)$ $=\left(\xi_{p}\Sigma_{p}(s)^{1/2}+\xi_{o}\Sigma_{o}(s)^{1/2}\right)^{2}$ (10)
where $\Sigma^{1/2}$ is the principal matrix square root of a positive semi-definite covariance matrix $\Sigma$. Figure  2 illustrates how the exploration policy gradually transitions toward the optimistic policy as a function of the evolving Wasserstein barycenter in one-dimensional action space in a fixed state.
Algorithm 1 summarizes the overall WBSAC procedure. During environment interaction, actions are sampled from the exploration policy $\pi_{e}$, which is dynamically formed as the Wasserstein barycenter of the pessimistic and optimistic policies using scheduled weights. In the training phase, the critic networks, both actor networks, and the SAC temperature parameter $\alpha$ are updated using transitions sampled from the replay buffer.
$\xi_{o}=0.1$
$\xi_{o}=0.9$
Figure 2: Exploration policy evolution. Initially, the exploration policy $\pi_{e}$ aligns with the pessimistic policy $\pi_{p}$. As learning proceeds, the exploration policy shifts toward the optimistic policy $\pi_{o}$.
Remark: Encouraging the agent to explore the parts of the action space that the two critics assign high value or disagree on their value after the exploitation phase, avoids the problem of unidirectional exploration in SAC, which was introduced by Ciosek et al. (2019) first. In other words, the inclusion of the standard deviation term (critic disagreement) in the optimistic actor’s objective explicitly encourages exploration towards regions of higher epistemic uncertainty, rather than uniformly sampling the upper bound of Q-value estimates.
Remark: By linking the exploration policy to the pessimistic policy at the beginning of the learning process, we prioritize exploitation and ensure that the transitions sampled from the exploration policy align with the probabilities expected under the pessimistic policy. As the training procedure proceeds, we encourage the agent to a small degree of exploration to escape the local optimum associated with lower-bound estimations (if any). The gradual interpolation from purely pessimistic to increasingly optimistic alleviates the phenomenon known as pessimistic underexploration, first discussed in Ciosek et al. (2019), in which an agent constrained by overly cautious value estimates can become trapped in suboptimal regions of the state–action space. WBSAC’s weighted-barycenter mechanism guides the policy beyond such a local optimum.
Proposition 1. For factorized Gaussian pessimistic and optimistic policies, the exploration policy $\pi_{e}$ (derived from Eq. 9 and 10) has its differential entropy, $H(\pi_{e}(s))$, lower-bounded for any state $s\in\mathcal{S}$ as:
$H\!\bigl{(}\pi_{e}(s)\bigr{)}\;\;\geq\;\;\xi_{p}\,H\!\bigl{(}\pi_{p}(s)\bigr{)}+\xi_{o}\,H\!\bigl{(}\pi_{o}(s)\bigr{)}.$ (11)
The proof is deferred to the Appendix. B.
Remark: Proposition 1 provides a theoretical basis for the exploration capability of $\pi_{e}$, which is essential for achieving broad coverage and collecting diverse data. As the weight $\xi_{o}$ for the optimistic policy increases, its contribution to this lower bound ensures $\pi_{e}$ maintains a capacity for broad action selection. This stochasticity promotes the generation of diverse data, as the agent samples a wider variety of actions, and leads to enhancing coverage of the state-action space.
Table 1: Average return over the last $10$ evaluations over 55 trials across MuJoCo environments.
Task WBSAC (ours) SAC DARC OAC
Ant-v5 $5564.8\pm 188.5$ $5306.5\pm 418.3$ $3926.2\pm 904.2$ $\mathbf{5867.0}\pm 292.3$
HalfCheetah-v5 $\mathbf{6466.4}\pm 1411.0$ $5409.7\pm 2219.1$ $4234.1\pm 1768.3$ $5210.5\pm 2353.8$
Walker2d-v5 $\mathbf{4417.0}\pm 703.0$ $3939.1\pm 262.0$ $3760.8\pm 485.3$ $4112.9\pm 217.4$
Humanoid-v5 $\mathbf{5701.2}\pm 463.4$ $5086.5\pm 170.0$ $4545.5\pm 780.8$ $4850.1\pm 279.1$
5 Experiments
We evaluate WBSAC performance on MuJoCo continuous control environments Todorov et al. (2012), DeepMind control suite tasks Tassa et al. (2018), and the PointMaze navigation task de Lazcano et al. (2023). We additionally provide a sensitivity analysis of hyperparameters. In addition to vanilla SAC, our comparative analysis includes DARC framework Lyu et al. (2022), which incorporates two actor networks to mitigate the overestimation bias by taking the minimum of value estimates from two actors and the underestimation bias by taking the maximum of the minimum value estimates from double critics. Our third baseline is OAC Ciosek et al. (2019), which is on top of SAC and focuses on directed exploration to enhance sample efficiency.
Performance on MuJoCo benchmarks. We evaluate WBSAC performance on four MuJoCo continuous control tasks Brockman (2016) via OpenAI Gym Brockman (2016), and provide averaged results over five seeds. In MuJoCo benchmark environments, the maximum number of interaction steps is $1000$ per episode. We evaluate the performance of the pessimistic policy every $5,000$ step. We use the moving average window of $40$ iterations.
Table 2: Average return over the last $10$ evaluations over 55 trials across DeepMind control suite tasks.
Domain-Task WBSAC (ours) DARC SAC OAC
Finger-Turn Easy 943.95 $\pm$ 17.55 937.73 $\pm$ 27.25 892.81 $\pm$ 85.34 937.47 $\pm$ 15.37
Finger-Turn Hard 910.96 $\pm$ 60.29 784.17 $\pm$ 138.09 856.08 $\pm$ 90.36 891.46 $\pm$ 39.44
Ball-in-cup Catch 983.93 $\pm$ 2.29 980.12 $\pm$ 2.67 982.60 $\pm$ 2.00 983.90 $\pm$ 2.09
Hopper Hop 130.14 $\pm$ 116.82 120.33 $\pm$ 75.24 116.33 $\pm$ 50.66 63.00 $\pm$ 49.23
Cheetah Run 740.16 $\pm$ 41.33 685.53 $\pm$ 75.57 647.55 $\pm$ 26.77 672.05 $\pm$ 38.82
Walker Walk 965.34 $\pm$ 6.31 852.31 $\pm$ 108.17 852.31 $\pm$ 115.41 965.80 $\pm$ 2.85
Walker Run 684.25 $\pm$ 18.10 527.50 $\pm$ 77.67 577.86 $\pm$ 59.49 671.08 $\pm$ 41.42
Humanoid Run 144.06 $\pm$ 11.55 1.23 $\pm$ 0.21 22.88 $\pm$ 36.82 98.77 $\pm$ 50.70
Figure 3 presents the average total episode reward and its standard deviation for five seeds. We observe that using two actors enhances the overall performance. As can be seen, our algorithm outperforms SAC in all four case studies. We emphasize that no hyperparameter tuning was performed on the Humanoid task. Table 1 provides the average return and standard deviation of the last ten evaluations. As observed, in all case studies except for Ant-v5, WBSAC achieves superior or comparable performance to OAC.
(a) Walker2d
(b) HalfCheetah
(c) Humanoid
(d) Ant
Figure 3: Performance comparison on MuJoCo environemnts. WBSAC outperforms SAC and DARC in all tasks and OAC in three.
(a) Ball in cup catch
(b) Cheetah-Run
(c) Finger turn-easy
(d) Finger turn-hard
(e) Hopper Hop
(f) Humanoid-Run
(g) Walker-Run
(h) Walker-Walk
Figure 4: Performance comparison on DeepMind control suite tasks. WBSAC consistently outperforms SAC and DARC, while it shows better or comparable performance with respect to OAC.
Performance on DeepMind control suite tasks. We further assess WBSAC performance on seven challenging tasks from DeepMind control suite Tassa et al. (2018) with both dense and sparse reward settings, which are known to be difficult for many off-policy, model-free reinforcement learning algorithms.
Figure 4 depicts the average return and its standard deviation for five seeds. The results demonstrate that WBSAC outperforms the baselines and successfully solves the challenging humanoid-run task on which other methods fail. Table 2 provides the average of last ten evaluation of the optimized policies.
Figure 5: Average coverage over 3 seeds in the PointMaze Medium-v3 navigation task with sparse reward.
Exploration capability. To assess the exploration capabilities of WBSAC, we conducted tests in a navigation maze, which is a modified version of the PointMaze Medium-v3 environment de Lazcano et al. (2023). In this PointMaze Medium-v3 setup, the agent’s objective is to guide a ball to an unknown goal location within the maze. The ball starts in the maze’s center, and we’ve designated two possible goal locations: the top-right and bottom-left corners. A sparse 010-1 reward is applied upon reaching the goal. We train the agent for $100\text{k}$ steps, with its policy being evaluated every $5000$ steps. We present the average return and average coverage across $3$ different random seeds.
(a) WBSAC
(b) SAC
(c) DARC
(d) OAC
(e) WBSAC
(f) SAC
(g) DARC
(h) OAC
Figure 6: State visitation heatmaps for pessimistic (first row) and exploration policies (second row).
Figure 5 indicates WBSAC achieves a higher coverage percentage than other baselines. Figure 6 presents exploration heatmaps for the PointMaze task from one representative seed to visually evaluate exploration efficacy in this sparse reward setting. As shown in Figure 6, WBSAC exploration policy achieves more comprehensive state-space coverage compared to baselines. In comparison to baselines, WBSAC wastes less resources on visiting non-rewarding regions of the maze, and focus its exploration on promising areas, which is because of directed exploration. The heatmap of our pessimistic policy shows that WBSAC successfully identifies multiple distinct paths to the goal, which indicates a more robust understanding of the task.
(a) Sensitivity to $\beta_{o}$.
(b) Sensitivity to $\lambda$.
Figure 7: Sensitivity analysis of WBSAC to hyperparameters in HalfCheetah-v5 domain.
Hyperparameter sensitivity. WBSAC relies on two hyperparameters, $\lambda$, which determines the transition rate from pessimistic strategy to optimistic behavior, and $\beta_{o}$, which controls the amount of uncertainty used to compute the optimistic policy. Figure 7(a) and Figure 7(b) demonstrate that the degree of optimism and the exploration schedule parameter affect the agent’s performance. For very low values of $\lambda$ the policy remains pessimistic for longer, which slows down exploration and convergence to optimal behavior. Conversely, very high values of $\lambda$ lead to a rapid shift towards the optimistic policy and can cause unstable learning or convergence to poor local optima. Finally, removing the influence of critic disagreement on the optimistic actor’s objective by setting the optimistic degree $\beta_{o}=0$ leads to a noticeable degradation in WBSAC’s performance.
6 Conclusion, limitations and future work
We present the WBSAC framework, a model-free deep reinforcement algorithm that uses dual actors. We use the Wasserstein barycenter of two pessimistic and optimistic policies to make a balance between exploration and exploitation. By adjusting the hyperparameters, WBSAC starts with a pessimistic strategy and gradually shifts to an optimistic behavior. This enhances sample efficiency and facilitates a smooth exploration process. This is validated through several experiments.
Despite its promising performance, the WBSAC framework presents a few limitations. Using a dual actor network setup leads to increased memory and computational cost compared to the state-of-the-art off-policy actor-critic algorithms. Furthermore, for policies represented by Gaussian distributions, the barycenter can often be derived in a straightforward, closed-form analytical solution. However, for more complex, non-Gaussian policy representations, the computation of the Wasserstein barycenter typically lacks a closed form and requires iterative numerical approaches. A future direction for this work is to compensate for this cost by focusing on the effective sampling experience replay buffer data, potentially with replay ratio scaling.
Appendix A Further discussion and broader impacts
The WBSAC exploration policy entropy lower bound ensures that the transition from pessimistic to optimistic behavior maintains a minimum level of action diversity. This enhanced state-action coverage has practical implications beyond immediate performance gains. For instance, an agent trained on a wider variety of experiences is less likely to overfit to a narrow subset of the environment. This enables better generalization to new tasks and enhances adaptability in unfamiliar environments, which are essential for effective real-world deployment. WBSAC’s mechanism of beginning with a pessimistic policy and gradually incorporating optimism could also offer a framework for safer exploration. The initial conservative behavior might help avoid catastrophic failures during early learning phases, especially in safety-critical systems.
Appendix B Proof of proposition 1
Proposition 1.
For factorized Gaussian pessimistic and optimistic policies, the exploration policy $\pi_{e}$ (derived from (9) and (10) has its differential entropy, $H(\pi_{e}(s))$, lower-bounded for any state $s\in\mathcal{S}$ as:
$H\!\bigl{(}\pi_{e}(s)\bigr{)}\;\;\geq\;\;\xi_{p}\,H\!\bigl{(}\pi_{p}(s)\bigr{)}+\xi_{o}\,H\!\bigl{(}\pi_{o}(s)\bigr{)}.$ (12)
Proof.
Consider $d$-dimensional Gaussian pessimistic policy $\pi_{p}(\cdot|s)=\mathcal{N}(\mu_{p}(s),\Sigma_{p}(s))$, and optimistic policy $\pi_{o}(\cdot|s)=\mathcal{N}(\mu_{o}(s),\Sigma_{o}(s))$, with $\Sigma_{p}(s)=\operatorname{diag}(\sigma_{p,1}^{2}(s),\ldots,\sigma_{p,d}^{2}(s)$, and $\Sigma_{o}(s)=\operatorname{diag}(\sigma_{o,1}^{2}(s),\ldots,\sigma_{o,d}^{2}(s))$, where $\sigma_{p,i}^{2}(s)$ and $\sigma_{o,i}^{2}(s)$ are the variances along dimension $i$. Considering the differential entropy definition as $H(\pi(s))=\frac{1}{2}\ln\det(2\pi e\Sigma(s))$ for $\pi(s)$, we start by pessimistic policy. Since $\Sigma_{p}(s)$ is diagonal, its square root is $\Sigma_{p}(s)^{1/2}=\operatorname{diag}(\sigma_{p,1}(s),\ldots,\sigma_{p,d}(s))$. Hence,
$\det\Sigma_{p}(s)=\prod_{i=1}^{d}(\sigma_{p,i}(s))^{2}$ (13)
and,
$\ln\det\Sigma_{p}(s)=2\sum_{i=1}^{d}\ln(\sigma_{p,i}(s))$ (14)
Therefore, the entropy becomes:
$H(\pi_{p}(s))=\frac{1}{2}\ln\det(2\pi e\Sigma_{p}(s))=\frac{1}{2}\ln(2\pi e)^{d}+\sum_{i=1}^{d}\ln\sigma_{p,i}(s)$ (15)
Similarly, $H(\pi_{o}(s))=\frac{1}{2}\ln(2\pi e)^{d}+\sum_{i=1}^{d}\ln\sigma_{o,i}(s)$. For the exploration policy $\pi_{e}(\cdot|s)=\mathcal{N}(\mu_{e}(s),\Sigma_{e}(s))$, with covariance $\Sigma_{e}(s)$ derived from (10), we have:
$\xi_{p}\Sigma_{p}(s)^{1/2}+\xi_{o}\Sigma_{o}(s)^{1/2}=\operatorname{diag}(\xi_{p}\sigma_{p,1}(s)+\xi_{o}\sigma_{o,1}(s),\ldots,\xi_{p}\sigma_{p,d}(s)+\xi_{o}\sigma_{o,d}(s))$ (16)
Therefore, $\Sigma_{e}(s)=\operatorname{diag}\left((\xi_{p}\sigma_{p,1}(s)+\xi_{o}\sigma_{o,1}(s))^{2},\ldots,(\xi_{p}\sigma_{p,d}(s)+\xi_{o}\sigma_{o,d}(s))^{2}\right)$ and the entropy $H(\pi_{e}(s))$ becomes:
$H(\pi_{e}(s))=\frac{1}{2}\ln(2\pi e)^{d}+\sum_{i=1}^{d}\ln(\xi_{p}\sigma_{p,i}(s)+\xi_{o}\sigma_{o,i}(s))$ (17)
To proof inequality 12, consider:
$\xi_{p}\sum_{i=1}^{d}\ln\sigma_{p,i}(s)+\xi_{o}\sum_{i=1}^{d}\ln\sigma_{o,i}(s)=\sum_{i=1}^{d}\left(\xi_{p}\ln\sigma_{p,i}(s)+\xi_{o}\ln\sigma_{o,i}(s)\right)$ (18)
Since the $\ln(x)$ is concave for $x>0$, Jensen’s inequality, for non-negative $\xi_{p}$, $\xi_{o}$ such that $\xi_{p}+\xi_{o}=1$, gives:
$\sum_{i=1}^{d}\left(\xi_{p}\ln\sigma_{p,i}(s)+\xi_{o}\ln\sigma_{o,i}(s)\right)\leq\sum_{i=1}^{d}\ln(\xi_{p}\sigma_{p,i}(s)+\xi_{o}\sigma_{o,i}(s))$ (19)
Because $\xi_{p}+\xi_{o}=1$, the constant term can be decomposed as:
$\frac{1}{2}\ln(2\pi e)^{d}=\xi_{p}\cdot\frac{1}{2}\ln(2\pi e)^{d}+\xi_{o}\cdot\frac{1}{2}\ln(2\pi e)^{d}$ (20)
Adding (20) to each side of (19), we obtain:
$\begin{split}&\xi_{p}{\color[rgb]{0,0,0}\underbrace{\left(\frac{1}{2}\ln\left((2\pi e)^{d}\right)+\sum_{i=1}^{d}\ln\sigma_{p,i}(s)\right)}_{{\color[rgb]{0,0,1}H(\pi_{p}(s))}}}+\xi_{o}{\color[rgb]{0,0,0}\underbrace{\left(\frac{1}{2}\ln\left((2\pi e)^{d}\right)+\sum_{i=1}^{d}\ln\sigma_{o,i}(s)\right)}_{{\color[rgb]{0,0,1}H(\pi_{o}(s))}}}\\ &\leq{\color[rgb]{0,0,0}\underbrace{\frac{1}{2}\ln\left((2\pi e)^{d}\right)+\sum_{i=1}^{d}\ln\left(\xi_{p}\sigma_{p,i}(s)+\xi_{o}\sigma_{o,i}(s)\right)}_{{\color[rgb]{0,0,1}H(\pi_{e})}}}\end{split}$ (21)
Hence,
$\xi_{p}H(\pi_{p}(s))+\xi_{o}H(\pi_{o}(s))\leq H(\pi_{e}(s))$ (22)
Appendix C Experimental details
For a fair comparison with baselines, we set the initial seeds of the computational packages and fix the seeds of the environments to ensure reproducibility of results given initial seed values. Additionally, instead of randomly sampling training and evaluation seeds, we force the seeds to come from disjoint sets Ciosek et al. [2019]. For WBSAC, all networks share the same architecture. We use two hidden layers of $256$ units, and the activation function is ReLU. All networks are trained with Adam optimizer Kingma and Ba [2014] with a learning rate of 3e-4. The hyperparameters of WBSAC are tuned specifically for the Ant-v5 case study using grid search over $\beta_{o}\in\{1,1.5,2,2.5,3,4\}$, $\lambda\in\{2,6,10,12,15,20\}$. Hyperparameters were kept constant across tasks. All the hyperparameters of SAC, DARC, and OAC baselines are reported in Table.  3. The results for our SAC baseline were generated using this publicly available PyTorch implementation from https://github.com/denisyarats/pytorch_sac. To ensure fair comparison, the hyperparameters for SAC baseline were configured as specified in the original SAC paper Haarnoja et al. [2018]. For the DARC baseline, we use the source code available at https://github.com/dmksjfl/SMR/blob/master/DARC.py, and for reproducing the OAC results we used https://github.com/microsoft/oac-explore. All experiments were conducted on NVIDIA A100-PCIE GPU with 40GB of RAM.
Table 3: Hyperparameters used across MuJoCo and DeepMind control suite Tasks.
Hyperparameter WBSAC SAC DARC OAC
Number of hidden layers 2 2 2 2
Number of hidden nodes 256 256 256 256
Activation ReLU ReLU ReLU ReLU
Batch size 256 256 256 256
Replay buffer size $10^{6}$ $10^{6}$ $10^{6}$ $10^{6}$
Discount factor 0.99 0.99 0.99 0.99
Target smoothing 0.005 0.005 0.005 0.005
Optimizer Adam Adam Adam Adam
Actor learning rate $3\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$
Critic learning rate $3\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$ $3\times 10^{-4}$
Entropy coefficient $0.2$ $0.2$ N/A $0.2$
Maximum log std 22 22 N/A 22
Minimum log std 20-20 20-20 N/A 20-20
Regularization parameter N/A N/A $0.005$ N/A
Noise clip N/A N/A $0.5$ N/A
Exploration noise N/A N/A $\mathcal{N}(0,0.1)$ N/A
Upper bound uncertainty coefficient $1.5$ N/A N/A 4.36
Exploration schedule $\lambda$ $10$ N/A N/A N/A
Shift multiplier $\sqrt{2\delta}$ N/A N/A N/A $3.69$
Appendix D Ablation study
Dynamic and fixed $\xi_{o}$. To evaluate the impact of our dynamic exploration mechanism, we provide a comparison between WBSAC using its variable weighting scheme for the Wasserstein barycenter and configurations that employ fixed weights. As can be seen clearly in Figure 8, our adaptive approach for varying the weights of the pessimistic and optimistic policies achieves better performance in comparison to configurations that bias the exploration policy to the pessimistic policy via a high $\xi_{p}$.
Figure 8: Ablation study on WBSAC performance under fixed and adaptive $\xi_{o}$.
Figure 9: Performance comparison of WBSAC and SAC with $4$ gradient steps per environment interaction on the Humanoid task.
Performance with more training steps. We further investigate the impact of increased update-to-data ratios (gradient steps per environment step). Figure 9 compares the performance of WBSAC and SAC with $4$ gradient steps per environment step, averaged over three seeds on the Humanoid task. WBSAC ultimately achieves a higher final average return than SAC under these experimental conditions.
Appendix E Code access
The source code for our experiments is included as a supplementary zip file.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our abstract and introduction sections present the paper’s contribution and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of the proposed work have been discussed in Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Complete proof of proposition 1 is presented in the Appendix. B.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Details for reproducibility, along with a link to the code, are provided.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: : We provide data open access with source code anonymous link in Appendix F.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Experimental setup and details are provided.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We provided the mean and standard deviation of the five trials’ results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Details of all compute resources are provided.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The work confirms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss our broader societal impacts in Appendix F.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: This work poses no high risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite all the used open-source baselines as references for acknowledgement of these assets.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide our code asset via an anonymized URL in Appendix F.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not interact with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not interact with human subjects in this work.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We used LLM for editing our manuscript.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.